# Learning Metrics that Maximise Power for Accelerated A/B-Tests

### v1

@@ 1. Introduction & Motivation
This paper synthesises, generalises and extends several of the aforementioned works into a general framework to learn A/B-testing metrics that maximise the statistical power they harness. We specifically extend the work of Kharitonov et al. (Kharitonov et al., 2017) to applications beyond web search, where the North Star can be delayed and insensitive. We highlight how their approach of maximising the average $z$-score does not accurately reflect downstream metric utility in our case, in that it does not penalise disagreement with the North Star sufficiently (i.e. type-III errors (Mosteller, 1948; Kaiser, 1960; Urbano et al., 2019)). Indeed: whilst this approach maximises the mean $z$-score, it does not necessarily improve the median $z$-score, and does not lead to improved statistical power in the form of reduced type-II error as a result.
Alternatively, optimising the learnt metric to minimise $p$-values —either directly or after applying a $\log$-transformation— more equitably divides gains over multiple experiments, leading to more statistically significant results instead of a few extremely significant results. Furthermore, we emphasise that learnt metrics are not meant to replace existing metrics, but rather to complement them. As such, their evaluation should be done through multiple hypothesis testing (with appropriate corrections (Shaffer, 1995)) if any of the North Star, available vetted proxies and surrogates, or learnt metrics are statistically significant under the considered treatment variant. We can then either adopt a conservative plug-in Bonferroni correction to temper type-I errors, or analyse synthetic A/A experiments to ensure the final procedure matches the required confidence level.
@@ 2.1. Statistical Significance Testing
We want to assess whether the mean of metric $m_{i}$ is statistically significantly higher under variant $A$ compared to variant $B$. To this end, we define a confidence level $\alpha$ (often $\alpha\approx 0.05$), corresponding to the false-positive-rate we deem acceptable. Then, we apply Welch’s $t$-test. The test statistic (also known as the $z$-score) for metric $m_{i}$ and the given variants is given by:
@@
(3) $p_{i}^{A\succ B}=2\cdot\min(\Phi(z_{i}^{A\succ B});1-\Phi(z_{i}^{A\succ B})).$
@@
When $p_{i}^{A\succ B}<\alpha$, we can confidently reject the null-hypothesis that $A$ and $B$ are equivalent w.r.t. the mean of metric $m_{i}$. Note that $z$-scores are signed, whereas two-tailed $p$-values are not. Indeed: $p_{i}^{A\succ B}\equiv p_{i}^{B\succ A}$, leaving room for faulty conclusions of directionality, known as type-III errors (Mosteller, 1948; Kaiser, 1960; Urbano et al., 2019). We discuss these later in depth.
A one-tailed $p$-value for the one-tailed null hypothesis $A\nsucc B$ is given by $1-\Phi(z_{i}^{A\succ B})$, and rejected when $<\frac{\alpha}{2}$. Throughout, we use two-tailed $p$-values unless mentioned otherwise.
@@ 2.1.1. $p$-value corrections
The above procedure is valid for a single metric, a single hypothesis, and importantly, a single decision. Nevertheless, this is not how experiments run in practice. Without explicit corrections on the $p$-values (or corresponding $z$-scores), violations of these assumptions lead to inflated false-positive-rates. We consider two common cases: a (conservative) multiple testing correction when an experiment has several treatments, and a sequential testing correction when experiments have no predetermined end-date or sample size at which to conclude. These corrections are applied as experiment-level corrections, to ensure that for any metric $m_{i}$ and variants $A,B$, the obtained $p$-values $p_{i}^{A\succ B}$ accurately reflect what they should reflect, yielding the specified coverage at varying confidence levels $\alpha$.
@@
We can equivalently apply this correction on $z$-scores instead, allowing us to directly compare $z$-scores across experiments with varying numbers of treatments. Recall that the percentile point function is the inverse of the CDF. We obtain a one-tailed $p$-value as $p=1-\Phi(z)$, and we reject the null hypothesis when $p<\frac{\alpha}{2}$. Now, instead, we reject when $p<\frac{\alpha}{2T}$. As such, computing corrected $z$-scores as $\bar{z}=z\frac{\Phi^{-1}(\frac{\alpha}{2})}{\Phi^{-1}(\frac{\alpha}{2T})}$ controls type-I errors effectively.
@@ 2.2. Learning Metrics that Maximise Sensitivity
These $z$-scores can be used exactly as before to obtain $p$-values. An intuitive property of the $z$-score, is that a relative $z$-score of $r=\frac{z^{A\succ B}_{\omega}}{z^{A\succ B}_{i}}$ implies that $\omega$ requires a factor $r^{2}$ fewer samples to achieve the same confidence level as $m_{i}$ (Chapelle et al., 2012). This can directly be translated to the cost of experimentation, as it allows us to run experiments for shorter time-periods or on smaller sub-populations.
@@
Here, $A\succ B$ implies that there is a known and vetted preference of variant $A$ over $B$ — typically because the North star or other guardrail metrics showed statistically significant improvements. We denote inconclusive experiments as $A~{}?~{}B$, implying statistically insignificant outcomes on the North Star. In rare cases, the North star might have gone up at the expense of important guardrail metrics, rendering conclusions ambiguous. Finally, $A\simeq B$ represents A/A experiments, where we know the null hypothesis to hold true (by design). The first set of experiments is used to measure type-III errors. Known and unknown outcomes are used to measure type-II errors, and A/A experiments can inform us about type-I errors. This dataset of past A/B experiments is collected and labelled by hand, from natural experiments occurring on the platform over time.
@@ 3.1. Learning Metrics that Maximise Power
When directly optimising $z$-scores, an implicit assumption is made that the utility we obtain from increased $z$-scores is linear. This is seldom a truthful characterisation of reality, considering how we wish to actually use these metrics downstream. We provide a toy example in Table 1, reporting $z$-scores and one-tailed $p$-values for two experiments and three possible metrics $m_{1},m_{2},m_{3}$, inspired by real data. In this toy example, we know that $A\succ B$, based on a hypothetical North Star metric. Nevertheless, as we do not know this beforehand, we typically test for the null hypothesis $A\simeq B$ with two-tailed $p$-values. In the table, this means that the outcome is statistically significant if the reported one-tailed values are $p<\frac{\alpha}{2}$ or $p>1-\frac{\alpha}{2}$. We report the power of every metric at varying confidence levels $\alpha\in\{0.05,0.01\}$, reporting whether (i) the null hypothesis is correctly rejected ($p<\frac{\alpha}{2}$, ), (ii) the outcome is inconclusive (?), i.e. a type-II error, or (iii) the null hypothesis is rejected, but for the wrong reason ($p>1-\frac{\alpha}{2}$, $\times$).
This latter case is deeply problematic, as it signifies disagreement with the North Star. Such errors have been described as type-III errors in the statistical literature (Mosteller, 1948; Kaiser, 1960; Urbano et al., 2019). Naturally, we would rather have a metric that fails to reject the null than one that confidently declares a faulty variant to be superior. Indeed, Deng and Shi argue that both directionality and sensitivity are desirable attributes for any metric (Deng and Shi, 2016). Nevertheless, considering candidate metrics in Table 1, type-III errors are not sufficiently penalised by the average $z$-score: metric $m_{3}$ maximises this objective despite yielding statistical power that is on par with a coin flip.
Directly maximising power might prove cumbersome, as it is essentially a discrete step-function w.r.t. the $z$-score, dependent on the confidence level $\alpha$. Instead, it comes natural to minimise the one-tailed $p$-value reported in Table 1. Indeed, the $p$-value transformation models diminishing returns for high $z$-scores, which allows type-III errors to be sufficiently penalised. When considering this objective, $m_{3}$ is clearly suboptimal whilst $m_{2}$ is preferred.
@@
Note that whilst direct optimisation of $p$-values is an improvement over myopic consideration of $z$-scores, there is another caveat: the “worst-case” loss of a type-III error is bounded at 1, which does not reflect our true utility function: metrics that disagree with the North Star are far less reliable than those that simply remain inconclusive. As such, we also consider another variant of the objective, where $\bar{p}=-p\log(1-p)$. Figure 1 provides visual intuition to clarify how this monotonic transformation on the $p$-values more heavily penalises type-III errors, whilst retaining the optimum. From a theoretical perspective, this function provides a convex relaxation for minimising the number of type-III errors a metric produces. As a result, we expect this surrogate to exhibit strong generalisation. We refer to this objective as minimising the $\log p$-value.
@@
Note that one could envision further extensions here where the confidence level $\alpha$ is directly incorporated into the objective function to maximise statistical power at a given confidence level $\alpha$. Nevertheless, we conjecture that their discrete nature might hamper effective optimisation and generalisation, compared to the strictly convex and smooth surrogate we obtain from the $\log p$-value.
@@ 4. Experiments & Discussion
Existing work on this topic used a private dataset from Yandex focused on web search experiments that ran between 2013–2014 (Kharitonov et al., 2017). They report type-I and -II errors for 8 metrics and a fixed 5% confidence level, over 118 A/B-tests and 472 A/A-tests.
In this work, we consider more general metrics that are relevant for use-cases beyond web search (i.a. user retention and various engagement signals). Furthermore, we report type-I/II/III errors at varying confidence levels, providing insights into the learnt metrics’ behaviour. For this, we leverage logs of past A/B-experiments on two large-scale short-video platforms with over 160 million monthly active users each: ShareChat and Moj. The datasets consist of 153 A/B-experiments total that ran in 2023, and over 25 000 A/A-pairs. In total, we have access to roughly 100 metrics detailing various types of interactions with the platform, engagements, and delayed retention signals. Because our dataset is limited in size (a natural consequence of the problem domain), we are bound to overfit when using all available metrics as input features. As such, we limit ourselves to 10 input metrics to learn from, and evaluate them w.r.t. the delayed North Star. This feature selection step also ensures that our linear model consists of fewer parameters, which increases practitioners’ and business stakeholders’ trust in its output. We focus on non-delayed signals, including activity metrics such as the number of sessions and active days, and counters for positive and negative feedback engagements of various types. These are selected through an analysis of their type-I/II/III errors w.r.t. the North Star, as well as their $z$-scores: focusing on metrics with high sensitivity and limited disagreement. The research questions we wish to answer empirically using this data, are the following:
@@
How do learnt metrics behave in terms of type-III errors?
@@ 4.1. Effectiveness of Learnt Metrics (RQ1)
Figure 3. Learnt metrics and their agreement with known North Star outcomes over varying confidence levels. A learnt metric that maximises the average $z$-score exhibits significant type-III error, which can be alleviated by minimising $p$-values instead.
@@ 4.2. Agreement with the North Star (RQ2)
From the obtained $z$-scores and $p$-values summarised in Table 2, we can additionally derive (dis-)agreement with the North Star, for varying confidence levels $\alpha$. We visualise these results in Figure 3: if the obtained $p$-value under a learnt metric is lower than $\alpha$, that metric yields a statistically significant result (agreement). If the obtained $p$-value for the alternative hypothesis (i.e. $B\succ A$ when we know $A\succ B$) is lower than $\alpha$, we have statistically significant disagreement, or type-III error. This is a capital sin we wish to avoid at all costs, as it severely diminishes the trust we can put in the learnt metric. If the $p$-value reveals a statistically insignificant result, we say the result is inconclusive, implying a type-II error. We observe that both the $z$-score-maximising metric and the heuristic approach fail to steer clear from type-III error. Optimising ($\log$)$p$-values instead, alleviates this issue. For this reason, we only consider these metrics for further evaluation. Indeed: an analysis of type-II error is rendered meaningless when type-III errors are present.
@@ 4.3. Power Increase from Learnt Metrics (RQ3–4)
Until now, we have leveraged experiments with known outcomes to assess sensitivity and agreement with the North Star. Now, we additionally consider A/A-experiments ($\mathcal{E}^{\simeq}$) and experiments with unknown outcomes ($\mathcal{E}^{?}$) to measure type-I and type-II error respectively. We measure this for the North Star, for the best-performing proxy metric that serves as input to the learnt metrics, and for learnt metrics that exhibit no empirical disagreement with the North Star. We plot the type-I error (i.e. fraction of A/A-pairs in $\mathcal{E}^{\simeq}$ that are statistically significant at confidence level $\alpha$) and the type-II error (i.e. fraction of A/B-pairs in {$\{\mathcal{E}^{+}\cup\mathcal{E}^{?}\}$ that are statistically insignificant at confidence level $\alpha$) for varying values of $\alpha$ in Figure 3(a). We observe that we are able to significantly reduce type-II errors compared to the North Star (up to 78%), whilst keeping type-I errors at the required level (i.e. $\alpha$). However, we also observe that the type-II error we obtain when using learnt metrics does not significantly improve over the top proxy metric, when considered in isolation.
Nonetheless, this is not how evaluation metrics are used in practice. Indeed, we track several metrics and can draw conclusions if any of them are statistically significant. As such, metrics should be evaluated on their complementary sensitivity. That is, we compute $p$-values for a set of metrics, apply a Bonferroni correction, and assess statistical significance. The statistical power that we obtain through this procedure is visualised in Figure 3(b). We consider either the North Star in isolation, the North Star in conjunction with the top-proxy, or a further combination with any learnt metric. Here, we observe that the learnt metric provides a substantial increase in statistical power: statistical power (i.e. $1-$ type-II error) is increased by up to 210% compared to the North Star alone, and 25–30% over the North Star plus proxies. Furthermore, as the Bonferroni correction is slightly conservative, we observe lower than expected type-I error for higher confidence levels $\alpha$. This implies that a more fine-grained multiple testing correction can further improve statistical power. We empirically observe that this works as expected, but its effects are negligible in practice.

### v2

@@ Abstract.
A/B-Testing; Evaluation Metrics; Statistical Power
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spaindoi: 10.1145/3637528.3671512isbn: 979-8-4007-0490-1/24/08ccs: General and reference Experimentationccs: Mathematics of computing Hypothesis testing and confidence interval computationccs: Computing methodologies Machine learning
@@ 1. Introduction & Motivation
This paper synthesises, generalises and extends several of the aforementioned works into a general framework to learn A/B-testing metrics that maximise the statistical power they harness. We specifically extend the work of Kharitonov et al. (Kharitonov et al., 2017) to applications beyond web search, where the North Star can be delayed and insensitive. We highlight how their approach of maximising the average $z$-score does not accurately reflect downstream metric utility in our case, in that it does not penalise disagreement with the North Star sufficiently (i.e. type-III/S errors (Mosteller, 1948; Kaiser, 1960; Gelman and Carlin, 2014; Urbano et al., 2019)). Indeed: whilst this approach maximises the mean $z$-score, it does not necessarily improve the median $z$-score, and does not lead to improved statistical power in the form of reduced type-II error as a result.
Alternatively, optimising the learnt metric to minimise $p$-values —either directly or after applying a $\log$-transformation— more equitably ditributes gains over multiple experiments, leading to more statistically significant results instead of a few extremely significant results. Furthermore, we emphasise that learnt metrics are not meant to replace existing metrics, but rather to complement them. As such, their evaluation should be done through multiple hypothesis testing (with appropriate corrections (Shaffer, 1995)) if any of the North Star, available vetted proxies and surrogates, or learnt metrics are statistically significant under the considered treatment variant. We can then either adopt a conservative plug-in Bonferroni correction to temper type-I errors, or analyse synthetic A/A experiments to ensure the final procedure matches the expected confidence level.
@@ 2.1. Statistical Significance Testing
We want to assess whether the mean of metric $m_{i}$ is statistically significantly higher under variant $A$ compared to variant $B$. To this end, we define a significance level $\alpha$ (often $\alpha\approx 0.05$), corresponding to the false-positive-rate we deem acceptable. Then, we apply Welch’s $t$-test. The test statistic (also known as the $z$-score) for metric $m_{i}$ and the given variants is given by:
@@
(3) $p_{i}^{A\neq B}=2\cdot\min(\Phi(z_{i}^{A\succ B});1-\Phi(z_{i}^{A\succ B})).$
@@
When $p_{i}^{A\neq B}<\alpha$, we can confidently reject the null-hypothesis that $A$ and $B$ are equivalent w.r.t. the mean of metric $m_{i}$. Note that $z$-scores are signed, whereas two-tailed $p$-values are not. Indeed: relabeling the variants changes the $z$-score but not the $p$-value, which leaves room for faulty conclusions of directionality, known as type-III errors (Mosteller, 1948; Kaiser, 1960; Urbano et al., 2019) or sign errors (Gelman and Carlin, 2014). We discuss these phenomena in detail, further in this article.
A one-tailed $p$-value for the one-tailed null hypothesis $A\nsucc B$ is given by $p_{i}^{A\nsucc B}=1-\Phi(z_{i}^{A\succ B})$, and rejected when $<\frac{\alpha}{2}$. Throughout, we use two-tailed $p$-values unless mentioned otherwise.
@@ 2.1.1. $p$-value corrections
The above procedure is valid for a single metric, a single hypothesis, and importantly, a single decision. Nevertheless, this is not how experiments run in practice. Without explicit corrections on the $p$-values (or corresponding $z$-scores), violations of these assumptions lead to inflated false-positive-rates. We consider two common cases: a (conservative) multiple testing correction when an experiment has several treatments, and a sequential testing correction when experiments have no predetermined end-date or sample size at which to conclude. These corrections are applied as experiment-level corrections, to ensure that for any metric $m_{i}$ and variants $A,B$, the obtained $p$-values accurately reflect what they should reflect, yielding the specified coverage at varying confidence levels $\alpha$.
@@
We can equivalently apply this correction on $z$-scores instead, allowing us to directly compare $z$-scores across experiments with varying numbers of treatments. Recall that the percentile point function is the inverse of the CDF. We obtain a one-tailed $p$-value as $p=1-\Phi(z)$, and we reject the one-tailed null hypothesis when $p<\frac{\alpha}{2}$. Now, instead, we reject when $p<\frac{\alpha}{2T}$. As such, computing corrected $z$-scores as $\bar{z}=z\frac{\Phi^{-1}(\frac{\alpha}{2})}{\Phi^{-1}(\frac{\alpha}{2T})}$ controls type-I errors effectively.
@@ 2.2. Learning Metrics that Maximise Sensitivity
These $z$-scores can be used exactly as before to obtain $p$-values. An intuitive property of the $z$-score, is that a relative $z$-score of $r=\frac{z^{A\succ B}_{\omega}}{z^{A\succ B}_{i}}$ implies that $\omega$ requires a factor $r^{2}$ fewer samples to achieve the same significance level as $m_{i}$ (Chapelle et al., 2012). This can directly be translated to the cost of experimentation, as it allows us to run experiments for shorter time-periods or on smaller sub-populations.
@@
Here, $A\succ B$ implies that there is a known and vetted preference of variant $A$ over $B$ — typically because the North star or other guardrail metrics showed statistically significant improvements. These experiments are further validated by replicating outcomes, observing long-term holdouts, or because the experiment was part of an intentional degradation test. We denote inconclusive experiments as $A~{}?~{}B$, implying statistically insignificant outcomes on the North Star. In rare cases, the North star might have gone up at the expense of important guardrail metrics, rendering conclusions ambiguous. We only include experiments into the inconclusive set for which we have a very strong intuition that something changed (and we “know” the null hypothesis should be rejected), but we are unable to make a confident directionality decision. This ensures that we can use this set to truly measure type-II errors. Finally, $A\simeq B$ represents A/A experiments, where we know the null hypothesis to hold true (by design). The first set of experiments is used to measure type-III/S errors. Known and unknown outcomes are used to measure type-II errors, and A/A experiments can inform us about type-I errors. This dataset of past A/B experiments is collected and labelled by hand, from natural experiments occurring on the platform over time.
@@ 3.1. Learning Metrics that Maximise Power
When directly optimising $z$-scores, an implicit assumption is made that the utility we obtain from increased $z$-scores is linear. This is seldom a truthful characterisation of reality, considering how we wish to actually use these metrics downstream. We provide a toy example in Table 1, reporting $z$-scores and one-tailed $p$-values for two experiments and three possible metrics $m_{1},m_{2},m_{3}$, inspired by real data. In this toy example, we know that $A\succ B$, based on a hypothetical North Star metric. Nevertheless, as we do not know this beforehand, we typically test for the null hypothesis $A\simeq B$ with two-tailed $p$-values. In the table, this means that the outcome is statistically significant if the reported one-tailed values are $p<\frac{\alpha}{2}$ or $p>1-\frac{\alpha}{2}$. We report the power of every metric at varying significance levels $\alpha\in\{0.05,0.01\}$, reporting whether (i) the null hypothesis is correctly rejected ($p<\frac{\alpha}{2}$, ✓ ), (ii) the outcome is inconclusive (?), i.e. a type-II error, or (iii) the null hypothesis is rejected, but for the wrong reason ($p>1-\frac{\alpha}{2}$, $\times$).
This latter case is deeply problematic, as it signifies disagreement with the North Star. Such errors have been described as type-III or type-S errors in the statistical literature (Mosteller, 1948; Kaiser, 1960; Gelman and Carlin, 2014; Urbano et al., 2019). Naturally, we would rather have a metric that fails to reject the null than one that confidently declares a faulty variant to be superior. Indeed, Deng and Shi argue that both directionality and sensitivity are desirable attributes for any metric (Deng and Shi, 2016). Nevertheless, considering candidate metrics in Table 1, type-III errors are not sufficiently penalised by the average $z$-score: metric $m_{3}$ maximises this objective despite yielding statistical power that is on par with a coin flip.
Directly maximising power might prove cumbersome, as it is essentially a discrete step-function w.r.t. the $z$-score, dependent on the significance level $\alpha$. Instead, it comes natural to minimise the one-tailed $p$-value reported in Table 1. Indeed, the $p$-value transformation models diminishing returns for high $z$-scores, which allows type-III errors to be sufficiently penalised. When considering this objective, $m_{3}$ is clearly suboptimal whilst $m_{2}$ is preferred.
@@
Note that whilst direct optimisation of $p$-values is an improvement over myopic consideration of $z$-scores, there is another caveat: the “worst-case” loss of a type-III/S error is bounded at 1, which does not reflect our true utility function: metrics that disagree with the North Star are far less reliable than those that simply remain inconclusive. As such, we also consider another variant of the objective, where $\bar{p}=-p\log(1-p)$. Figure 1 provides visual intuition to clarify how this monotonic transformation on the $p$-values more heavily penalises type-III/S errors, whilst retaining the optimum. From a theoretical perspective, this function provides a convex relaxation for minimising the number of type-III/S errors a metric produces. As a result, we expect this surrogate to exhibit strong generalisation. We refer to this objective as minimising the $\log p$-value.
@@
Note that one could envision further extensions here where the significance level $\alpha$ is directly incorporated into the objective function to maximise statistical power at a given significance level $\alpha$. Nevertheless, we conjecture that their discrete nature might hamper effective optimisation and generalisation, compared to the strictly convex and smooth surrogate we obtain from the $\log p$-value.
@@ 4. Experiments & Discussion
Existing work on this topic used a private dataset from Yandex focused on web search experiments that ran between 2013–2014 (Kharitonov et al., 2017). They report type-I and -II errors for 8 metrics and a fixed 5% significance level, over 118 A/B-tests and 472 A/A-tests.
In this work, we consider more general metrics that are relevant for use-cases beyond web search (i.a. user retention and various engagement signals). Furthermore, we report type-I/II/III/S errors at varying significance levels, providing insights into the learnt metrics’ behaviour. For this, we leverage logs of past A/B-experiments on two large-scale short-video platforms with over 160 million monthly active users each: ShareChat and Moj. The datasets consist of 153 A/B-experiments (of which 58 were conclusive) total that ran in 2023, and over 25 000 A/A-pairs. In total, we have access to roughly 100 metrics detailing various types of interactions with the platform, engagements, and delayed retention signals. Because our dataset is limited in size (a natural consequence of the problem domain), we are bound to overfit when using all available metrics as input features. As such, we limit ourselves to 10 input metrics to learn from, and evaluate them w.r.t. the delayed North Star. This feature selection step also ensures that our linear model consists of fewer parameters, which increases practitioners’ and business stakeholders’ trust in its output. We focus on non-delayed signals, including activity metrics such as the number of sessions and active days, and counters for positive and negative feedback engagements of various types. These are selected through an analysis of their type-I/II/III/S errors w.r.t. the North Star, as well as their $z$-scores: focusing on metrics with high sensitivity and limited disagreement. The research questions we wish to answer empirically using this data, are the following:
@@
How do learnt metrics behave in terms of type-III/S errors?
@@ 4.1. Effectiveness of Learnt Metrics (RQ1)
Figure 3. Learnt metrics and their agreement with known North Star outcomes over varying significance levels. A learnt metric that maximises the average $z$-score exhibits significant type-III/S error, which can be alleviated by minimising $p$-values instead.
@@ 4.2. Agreement with the North Star (RQ2)
From the obtained $z$-scores and $p$-values summarised in Table 2, we can additionally derive (dis-)agreement with the North Star, for varying significance levels $\alpha$. We visualise these results in Figure 3: if the obtained $p$-value under a learnt metric is lower than $\alpha$, that metric yields a statistically significant result (agreement). If the obtained $p$-value for the alternative hypothesis (i.e. $B\succ A$ when we know $A\succ B$) is lower than $\alpha$, we have statistically significant disagreement, or type-III error. This is a capital sin we wish to avoid at all costs, as it severely diminishes the trust we can put in the learnt metric. If the $p$-value reveals a statistically insignificant result, we say the result is inconclusive, implying a type-II error. We observe that both the $z$-score-maximising metric and the heuristic approach fail to steer clear from type-III error. Optimising ($\log$)$p$-values instead, alleviates this issue. For this reason, we only consider these metrics for further evaluation. Indeed: an analysis of type-II error is rendered meaningless when type-III errors are present.
@@ 4.3. Power Increase from Learnt Metrics (RQ3–4)
Until now, we have leveraged experiments with known outcomes to assess sensitivity and agreement with the North Star. Now, we additionally consider A/A-experiments ($\mathcal{E}^{\simeq}$) and experiments with unknown outcomes ($\mathcal{E}^{?}$) to measure type-I and type-II error respectively. We measure this for the North Star, for the best-performing proxy metric that serves as input to the learnt metrics, and for learnt metrics that exhibit no empirical disagreement with the North Star. We plot the type-I error (i.e. fraction of A/A-pairs in $\mathcal{E}^{\simeq}$ that are statistically significant at significance level $\alpha$) and the type-II error (i.e. fraction of A/B-pairs in {$\{\mathcal{E}^{+}\cup\mathcal{E}^{?}\}$ that are statistically insignificant at significance level $\alpha$) for varying values of $\alpha$ in Figure 4(a). We observe that we are able to significantly reduce type-II errors compared to the North Star (up to 78%), whilst keeping type-I errors at the required level (i.e. $\alpha$). However, we also observe that the type-II error we obtain when using learnt metrics does not significantly improve over the top proxy metric, when considered in isolation.
Nonetheless, this is not how evaluation metrics are used in practice. Indeed, we track several metrics and can draw conclusions if any of them are statistically significant. As such, metrics should be evaluated on their complementary sensitivity. That is, we compute $p$-values for a set of metrics, apply a Bonferroni correction, and assess statistical significance. The statistical power that we obtain through this procedure is visualised in Figure 4(b). We consider either the North Star in isolation, the North Star in conjunction with the top-proxy, or a further combination with any learnt metric. Here, we observe that the learnt metric provides a substantial increase in statistical power: statistical power (i.e. $1-$ type-II error) is increased by up to a relative 210% compared to the North Star alone, and 25–30% over the North Star plus proxies. Furthermore, as the Bonferroni correction is slightly conservative, we observe lower than expected type-I error for higher significance levels $\alpha$. This implies that a more fine-grained multiple testing correction can further improve statistical power. We empirically observe that this works as expected, but its effects are negligible in practice.